OpenStudy (lena772):

Which of the following statements best describes the graph of –5x + 2y = 1? It is a curve joining the points (–5, 2), (2, 3), and (4, 1). It is a curve joining the points (–1, –3), (–1, –3), and (1, 5). It is a straight line joining the points (1, 3), (3, 8), and (–3, –7). It is a straight line joining the points (4, –3), (–1, 2), and (–4, 5).

OpenStudy (ranga):

We know –5x + 2y = 1 is an equation of a straight line. Try the given points to see if they satisfy the equation.

OpenStudy (lena772):

-5+6=1 -15+16=1 15+-14=1

OpenStudy (lena772):

C

OpenStudy (ranga):

good! yes!